OpenStudy (anonymous):

*** Will give you a medal *** A circle is inscribed in a square. Write and simplify an expression for the ratio of the area of the square to the area of the circle. For a circle inscribed in a square, the diameter of the circle is equal to the side length of the square.

OpenStudy (anonymous):

The required ratio is D^2/(PI D^2/4) = 4/[\Pi \]

OpenStudy (anonymous):

so how would i figure that out ?

OpenStudy (dumbcow):

let side of square be "x" , diameter is "x" area of square = x^2 area of circle = pi*(x/2)^2 \[ratio = \frac{x^{2}}{\pi x^{2}/4} = \frac{4}{\pi}\]

OpenStudy (anonymous):

is that the answer ?

OpenStudy (anonymous):



OpenStudy (anonymous):

@dumbcow is that the answer ?

OpenStudy (dumbcow):

yes